# Probabilistic Analysis for Scalogram Ridges in Signal Processing

### v2

@@ 4 Definition of ridges
We start the definition of ridges by recalling existing results when the input signal is noise-free. With AHM and the slowly varying condition specified in (1), we have the following approximation for the AWT of $f$ fulfilling the AHM [2011synchrosqueezed, Estimate 3.5]:
@@
To analyze the ridges in the scalogram of $f$, we make the following observations. First, if the amplitude of the $m$-th IMT function in $f$, where $m\in\{1,\ldots,M\}$, is very small, its contribution to the scalogram may be buried by spectral leakage caused by the time-varying nature of the frequencies and amplitudes. This can make detecting the $m$-th IF, $\phi^{\prime}_{m}$, from the scalogram challenging. Conversely, if one component has a very large amplitude, its spectral leakage may dominate the scalogram, potentially masking the contributions of other components. Second, if the spectral energy of the mother wavelet $\psi$ is spread over a broad frequency range rather than being tightly concentrated, the components in (14) may overlap significantly, making it harder to isolate individual components or accurately infer the IFs. To build intuition about ridge definitions for noise-free AMH, we focus on the scenario where the impact of spectral leakage in the scalogram of the noise-free AHM is negligible. This holds, for example, when the support of $\hat{\psi}$ is concentrated around [1B,1+B][1-B,1+B], where $0<B<\delta/(1+\delta)$ and $\delta$ is the constant mentioned in the spectral separation condition (2). Under this assumption, (14) can be further approximated as follows:
$\left|W_{f}(t,s)\right|\approx\frac{1}{2}\left|\sum_{m=1}^{M}A_{m}(t)e^{i2\pi\phi_{m}(t)}\widehat{\psi}(2\pi s\phi^{\prime}_{m}(t))\right|=\frac{1}{2}\overset{M}{\underset{m=1}{\sum}}A_{m}(t)\left|\widehat{\psi}\left(2\pi s\phi^{\prime}_{m}(t)\right)\right|.$ (15)
@@
which is a singleton under Assumption 2. By (15), for any $t\in\mathbb{R}$, the local maxima of $\left|W_{f}(t,\cdot)\right|$ approximately occur at $\frac{\omega_{\psi}}{\phi^{\prime}_{1}(t)},\frac{\omega_{\psi}}{\phi^{\prime}_{2}(t)},\ldots,\frac{\omega_{\psi}}{\phi^{\prime}_{M}(t)}$, which are inversely related to the IFs $\{\phi^{\prime}_{1}(t),\phi^{\prime}_{2}(t),\ldots,\phi^{\prime}_{M}(t)\}$. The observation above leads to the following definition of ridges [lilly2010analytic, Definitions 2.1 and 2.2] (see also [meignen2023new, Section B]).
@@
that is, the ridge captures the IF information. See Figure 1 for an example of the graph of $s_{f}$ and its relationship with the IF $\phi^{\prime}(t)$. Based on the proof technique using the implicit function theorem presented in [colominas2023iterative], the functions $s_{f}(t)$ is continuously differentiable and hence constitutes a ridge. Note that [colominas2023iterative] analyzes the ridge of the spectrogram generated by the short-time Fourier transform. Since a similar proof can be carried out to the scalogram, we skip the details here.
When $f$ satisfies the AHM (1) with $M\geq 2$, (17) might be problematic. First, due to the time-varying amplitude, $s_{f}(t)$ might alternately capture the IF of different IMT functions. For example, consider the case where $M=2$, $\phi_{m}^{\prime}(t)=\xi_{m}>0$ are constants, and $\xi_{2}>2\xi_{1}$. If $A_{1}(t)>2A_{2}(t)$ for t<1t<-1 and $A_{2}(t)>2A_{1}(t)$ for $t>1$, then $s_{f}(t)\approx\omega_{\psi}/\phi^{\prime}_{1}(t)$ when t<1t<-1 and $s_{f}(t)\approx\omega_{\psi}/\phi^{\prime}_{2}(t)$ when $t>1$. On the other hand, note that when the spectral leakage from other IMT functions of $f$ is smaller than $A_{m}(t)|\widehat{\psi}(2\pi s\phi^{\prime}_{m}(t))|$, $s_{f,m}(t)$ is close to $\phi^{\prime}_{m}(t)$ and we have $M$ local maxima. In this case, to utilize the definition of ridges, consider
@@
where $\underline{b}_{1},\overline{b}_{1},\underline{b}_{2},\overline{b}_{2},\ldots,\underline{b}_{M},\overline{b}_{M}$ are continuous functions defined on $\mathbb{R}$ satisfying
@@
Similarly, by the same proof [colominas2023iterative] we can prove that $s_{f,m}(t)$ are continuously differentiable so they are ridge curves by definition. Hereafter, we use the graphs of $s_{f}$ and $\{s_{f,m}\}_{m=1,\ldots,M}$ to represent the ridge curves of the noise-free function satisfying the AHM and call them the ridge curves to simplify the nomination.
@@
In practice, information about the IFs of $f$, and hence $B_{m}(t)$, may sometimes be available through prior knowledge. However, more often than not, this information is unavailable and must be estimated. For an example of an algorithm addressing this estimation challenge, see [Laurent_Meignen_2021]. This algorithm, jointly with its theoretical analysis, fall in Steps III and IV, and lie beyond the scope of this paper. From a broader signal processing perspective, once this definition is established, the next step is to design an appropriate algorithm to extract the ridge curves $\{s_{f,m}\}_{m=1,\ldots,M}$ from the scalogram.
@@ 5.1 Local properties of ridge curves
Suppose $f\in L^{2}$, $\Phi$ a stationary Gaussian process satisfying Assumptions 1 and 3, and $\psi$ a mother wavelet satisfying Assumption 2. Then, we have
@@
Suppose $f\in L^{2}$, $\Phi$ a stationary Gaussian process satisfying Assumptions 1 and 3, and $\psi$ a mother wavelet satisfying Assumption 2. With probability one, the set-valued function $s_{Y}$ is upper hemicontinuous on $\mathbb{R}$. That is, almost surely, for any $t\in\mathbb{R}$ and any open set $U$ with $s_{Y}(t)\subset U$, there exists $\delta>0$ such that $s_{Y}(t^{\prime})\subset U$ for any $t^{\prime}\in(t-\delta,t+\delta)$. Moreover, if $f$ is an adapted harmonic signal as defined in (1) with $M\geq 1$, then almost surely, the set-valued functions $s_{Y,1},\ldots,s_{Y,M}$, defined in (20), are also upper hemicontinuous on $\mathbb{R}$.
The upper hemicontinuity of $s_{Y}$ implies that as $t^{\prime}$ approaches $t$, the maximizers of $S_{Y}(t^{\prime},\cdot)$ get arbitrarily close to some of the maximizers of $S_{Y}(t,\cdot)$, as illustrated in Figure 4. The proof of Theorem 3 is provided in Section 7.5, which relies on Berge’s maximum theorem [hu1997handbook, Theorem 3.4] (see also [aliprantis2006infinite, p. 570] and Lemma 5 in the appendix).
Theorems 2 and 3 imply that, with probability one, the restriction of $s_{Y}$ on the dense set $\mathbb{Q}$, denoted by $s_{Y}|_{\mathbb{Q}}$, is continuous. By [MR1039321, Lemma (4.3.16)] (see also Lemma 6 in the appendix), $s_{Y}|_{\mathbb{Q}}$ is extendable to a continuous function defined on a $G_{\delta}$-set containing $\mathbb{Q}$. Here, a $G_{\delta}$-set refers to a subset of $\mathbb{R}$ that is a countable intersection of open sets. Next, we provide more insights into this $G_{\delta}$-set.
@@ 5.2 Deviation of ridge points caused by noise
$\mathbb{P}\left(s_{Y}(t)\in I\right)\geq 1-\textup{exp}\left(-\frac{(\triangle_{I}-\mu_{I})^{2}}{\sigma_{I}^{2}}\right)$ (24)
@@
The proof of Theorem 4 is presented in Section 7.7. To prove Theorem 4 and derive an upper bound for the term $\mathbb{E}[\underset{s\in I^{c}}{\max}|W_{\Phi}(t,s)|]$ in (22), we need to extend the classical Gaussian concentration inequality for Lipschitz functions of real-valued Gaussian elements [kuhn2023maximal], Dudley’s theorem, and the Borell-TIS inequality [talagrand2014upper, vershynin2018high], to account for complex-valued Gaussian elements arising from the AWT of $\Phi$. To keep the focus on the discussion of ridge deviation, these extensions are provided in Section 5.4.
@@
The bound is expressed in terms of the threshold $\varepsilon$ and the following ratios:
$\frac{\mu_{1,m}}{L_{m}},\frac{\mu_{2,m}}{L_{m}},\frac{\mu_{3,m}^{2}}{L_{m}},\frac{\mu_{4,m}^{2}}{L_{m}},\frac{\sigma_{m,1}^{2}}{L^{2}_{m}},\frac{\sigma_{m,2}^{2}}{L^{2}_{m}},\frac{\sigma_{m,3}^{2}}{L_{m}},\frac{\sigma_{m,4}^{2}}{L_{m}},$
which are small when the noise strength is low relative to the clean signal; that is, the error control is related to the signal-to-noise ratio (SNR). The proof of Theorem 5 is presented in Section 7.8.
@@ 5.3 Numerical simulations
The result is shown in Figure 6. The left panel displays a scatter plot of all $10^{4}$ trials, showing the relationship between the SNR (in dB) and $\triangle$ or $\widetilde{\triangle}$. The right panel presents a box plot summarizing the ridge deviations within different SNR intervals. The result shows that when the SNR is relatively high (resp. low), the ridge point $s_{Y,m}(t)$ of the noisy signal $Y$ is unlikely (resp. likely) to deviate too far from the ridge point $s_{f,m}(t)$ of the clean signal $f$. This numerical result echos results shown in Theorem 5, providing a bound on the probability that the deviation $|s_{Y,m}(t)-s_{f,m}(t)|$ exceeds a given threshold.
@@ 5.4 Dudley’s theorem and the Borell-TIS inequality for AWT
To derive Theorems 4 and 5, a key challenge is estimating the tail probabilities of the following random variables arising from the AWT of the Gaussian process $\Phi$:
@@
Our estimation method is inspired by the Borell-TIS inequality [talagrand2014upper, vershynin2018high], which provides a tail probability estimate for the supremum of a real-valued non-stationary Gaussian process over an index set. In the proof of the Borell-TIS inequality, the Gaussian concentration inequality for Lipschitz functions of real-valued Gaussian random vectors [kuhn2023maximal] plays an important role. Lemma 1 below generalizes the concentration inequality in [kuhn2023maximal] to accommodate complex-valued and circularly symmetric Gaussian random vectors, whose proof is provided in Section 7.9.
@@
Note that in cases (31) and (32), the supremum is taken over the continuous index sets $I$ and $I^{c}$, respectively. In order to apply the dominated convergence theorem to extend the result in Lemma 1 to the continuous index set case, we need to ensure the convergence of the expected value of the supremum. Details of this extension are provided in the proof of Theorem 4. In [dudley1967sizes, talagrand2014upper], Dudley’s theorem provides an upper bound for the expected maximum of a centered real-valued Gaussian field over a compact set. Because the AWT of the Gaussian process $\Phi$ is a complex-valued random process and the index set $I^{c}$ for case (31) is noncompact, we present a complex-valued version of Dudley’s theorem for the expectation of the supremum of the process $|W_{\Phi}(t,\cdot)|$ over the noncompact set $0,\infty)$ in Lemma 3. The statement and proof of Lemma 3 rely on estimates of $\mathbb{E}[|W_{\Phi}(0,s)|^{2}]$ for large scale $s$ and the distance function $d_{W_{\Phi}}$ as follows.
@@
In [dudley1967sizes, talagrand2014upper], the expectation of the maximum of a centered real-valued Gaussian field over a compact set is bounded above by the Dudley entropy integral. Lemma 3 extends this bound, providing an exact upper bound for the expectation of the maximum of the modulus of a circularly symmetric and nonstationary complex Gaussian process over $(0,\infty)$. The upper bound given in (37) increases as the long-memory parameter $H$ and the decay rate parameter $\gamma$ of the spectral density function of $\Phi$ decrease to zero. This relationship echoes Slepian’s inequality and the Sudakov-Fernique inequality [vershynin2018high, Chapter 7], which state that, for two real-valued Gaussian processes with the same mean, the process with higher correlation exhibits a greater expected maximum. The proof of Lemma 3 is provided in Section 7.11. Finally, to estimate the tail probabilities of the random variables in (33), we introduce the complex-valued version of the Borell-TIS inequality as follows.
@@
The proof of Lemma 4 is provided in Section 7.12. Last but not least, our work bridges and extends classical maximum theorems and maximum inequalities to address a key question in signal processing: how noise perturbs the ridge curves of adapted harmonic signals that encode the IF information.

### v3

@@ 4 Definition of ridges
We start the definition of ridges by recalling existing results when the input signal is noise-free. With AHM and the slowly varying condition specified in (1), we have the following approximation for the AWT of $f$ fulfilling the AHM [10, Estimate 3.5]:
@@
To analyze the ridges in the scalogram of $f$, we make the following observations. First, if the amplitude of the $m$-th IMT function in $f$, where $m\in\{1,\ldots,M\}$, is very small, its contribution to the scalogram may be buried by spectral leakage caused by the time-varying nature of the frequencies and amplitudes. This can make detecting the $m$-th IF, $\phi^{\prime}_{m}$, from the scalogram challenging. Conversely, if one component has a very large amplitude, its spectral leakage may dominate the scalogram, potentially masking the contributions of other components. Second, if the spectral energy of the mother wavelet $\psi$ is spread over a broad frequency range rather than being tightly concentrated, the components in (14) may overlap significantly, making it harder to isolate individual components or accurately infer the IFs. To build intuition about ridge definitions for noise-free AMH, we focus on the scenario where the impact of spectral leakage in the scalogram of the noise-free AHM is negligible. This holds, for example, when the support of $\hat{\psi}$ is concentrated around [1B,1+B][1-B,1+B], where $0<B<\delta/(1+\delta)$ and $\delta$ is the constant mentioned in the spectral separation condition (2). Under this assumption, the regions where the values of the functions $\{s\mapsto|\widehat{\psi}\left(2\pi s\phi^{\prime}_{m}(t)\right)|\}_{m=1,2,\ldots,M}$ are non-negligible are nearly non-overlapping. Consequently, (14) can be further approximated as follows:
$\left|W_{f}(t,s)\right|\approx\frac{1}{2}\overset{M}{\underset{m=1}{\sum}}A_{m}(t)\left|\widehat{\psi}\left(2\pi s\phi^{\prime}_{m}(t)\right)\right|.$ (15)
@@
which is a singleton under Assumption 2. By (15), for any $t\in\mathbb{R}$, the local maxima of $\left|W_{f}(t,\cdot)\right|$ approximately occur at $\frac{\omega_{\psi}}{\phi^{\prime}_{1}(t)},\frac{\omega_{\psi}}{\phi^{\prime}_{2}(t)},\ldots,\frac{\omega_{\psi}}{\phi^{\prime}_{M}(t)}$, which are inversely related to the IFs $\{\phi^{\prime}_{1}(t),\phi^{\prime}_{2}(t),\ldots,\phi^{\prime}_{M}(t)\}$. The observation above leads to the following definition of ridges [26, Definitions 2.1 and 2.2] (see also [31, Section B]).
@@
that is, the ridge captures the IF information. See Figure 1 for an example of the graph of $s_{f}$ and its relationship with the IF $\phi^{\prime}(t)$. Based on the proof technique using the implicit function theorem presented in [8], the functions $s_{f}(t)$ is continuously differentiable and hence constitutes a ridge. Note that [8] analyzes the ridge of the spectrogram generated by the short-time Fourier transform. Since a similar proof can be carried out to the scalogram, we skip the details here.
When $f$ satisfies the AHM (1) with $M\geq 2$, (17) might be problematic. First, due to the time-varying amplitude, $s_{f}(t)$ might alternately capture the IF of different IMT functions. For example, consider the case where $M=2$, $\phi_{m}^{\prime}(t)=\xi_{m}>0$ are constants, and $\xi_{2}>2\xi_{1}$. If $A_{1}(t)>2A_{2}(t)$ for t<1t<-1 and $A_{2}(t)>2A_{1}(t)$ for $t>1$, then $s_{f}(t)\approx\omega_{\psi}/\phi^{\prime}_{1}(t)$ when t<1t<-1 and $s_{f}(t)\approx\omega_{\psi}/\phi^{\prime}_{2}(t)$ when $t>1$. On the other hand, when the spectral leakage from other IMT functions of $f$ is smaller than $A_{m}(t)|\widehat{\psi}(2\pi s\phi^{\prime}_{m}(t))|$, the cross-section of the scalogram at time $t$ exhibits $M$ local maxima. In this case, to utilize the definition of ridges, consider
@@
where $\underline{b}_{1},\overline{b}_{1},\underline{b}_{2},\overline{b}_{2},\ldots,\underline{b}_{M},\overline{b}_{M}$ are continuous on $\mathbb{R}$ and satisfy
@@
Similarly, by the same argument as in [8], we can prove that $s_{f,m}(t)$ are continuously differentiable, and hence they are ridge curves by definition. Hereafter, we use the graphs of $s_{f}$ and $\{s_{f,m}\}_{m=1,\ldots,M}$ to represent the ridge curves of the noise-free function satisfying the AHM, and we refer to them collectively as the ridge curves for simplicity.
@@
In this work, we focus on the local properties of ridge curves of the noisy signal $Y$, which will be defined in Section 4.1, and on their deviation from those of the noise-free signal $f$. Accordingly, we suppose that the scalogram $S_{f}(t,\cdot)$ of the noise-free signal exhibits $M$ well-separated local maxima at each time $t$ with known locations. This assumption ensures the existence of the interval functions $\underline{b}_{m}$ and $\overline{b}_{m}$.
The choice of $\underline{b}_{m}$ and $\overline{b}_{m}$ is not unique. When additional information about $f$ is available, the width of the ring-shaped regions $\{(t,s)\mid\underline{b}_{m}(t)\leq s\leq\overline{b}_{m}(t)\}_{m=1}^{M}$ may be reduced. In practice, the IFs of $f$, and hence the intervals $B_{m}(t)$, must be estimated. An example of an algorithm addressing this estimation problem is given in [24]. This algorithm, jointly with its theoretical analysis, falls in Steps III and IV, and lies beyond the scope of this paper.
Conditions ensuring the existence of well-separated local maxima in time-frequency representations have been studied for specific signal classes, such as purely harmonic modes or parallel linear chirps, primarily in the context of the spectrogram; see, for example, [32]. Extending such results to the situation that two IMT functions have close IFs, or even crossover IFs, constitutes an important but challenging problem. Handling this challenge is related to analyzing the spectral interference behavior of CWT. It is also important to study the case when the IF or AM fail the slowly varying condition. Since these challenges are out of the scope of this paper, we leave them to our future work.
@@ 5.1 Local properties of ridge curves
Suppose $f\in L^{2}$, $\Phi$ is a stationary Gaussian process satisfying Assumptions 1 and 3, and $\psi$ is a mother wavelet satisfying Assumption 2. Then, we have
@@
Suppose $f\in L^{2}$, $\Phi$ is a stationary Gaussian process satisfying Assumptions 1 and 3, and $\psi$ is a mother wavelet satisfying Assumption 2. With probability one, the set-valued function $s_{Y}$ is upper hemicontinuous on $\mathbb{R}$. That is, almost surely, for any $t\in\mathbb{R}$ and any open set $U$ with $s_{Y}(t)\subset U$, there exists $\delta>0$ such that $s_{Y}(t^{\prime})\subset U$ for any $t^{\prime}\in(t-\delta,t+\delta)$. Moreover, if $f$ is an adaptive harmonic signal as defined in (1) with $M\geq 1$, then almost surely, the set-valued functions $s_{Y,1},\ldots,s_{Y,M}$, defined in (20), are also upper hemicontinuous on $\mathbb{R}$.
The upper hemicontinuity of $s_{Y}$ implies that as $t^{\prime}$ approaches $t$, the maximizers of $S_{Y}(t^{\prime},\cdot)$ get arbitrarily close to some of the maximizers of $S_{Y}(t,\cdot)$, as illustrated in Figure 4. The proof of Theorem 2 is provided in Section 7.5, which relies on Berge’s maximum theorem [21, Theorem 3.4] (see also [2, p. 570] and Lemma 5 in the appendix).
Theorems 1 and 2 imply that, with probability one, the restriction of $s_{Y}$ on the dense set $\mathbb{Q}$, denoted by $s_{Y}|_{\mathbb{Q}}$, is continuous. By [17, Lemma (4.3.16)] (see also Lemma 6 in the appendix), $s_{Y}|_{\mathbb{Q}}$ is extendable to a continuous function defined on a $G_{\delta}$-set containing $\mathbb{Q}$. Here, a $G_{\delta}$-set refers to a subset of $\mathbb{R}$ that is a countable intersection of open sets. Next, we provide more insights into this $G_{\delta}$-set.
@@ 5.2 Deviation of ridge points caused by noise
$\mathbb{P}\left(s_{Y}(t)\in I\right)\geq 1-\textup{exp}\left(-\frac{(\triangle_{I}-\mu_{I})^{2}}{\sigma_{I}^{2}}\right),$ (24)
@@
The proof of Theorem 3 is presented in Section 7.7. To prove Theorem 3 and derive an upper bound for the term $\mathbb{E}[\underset{s\in I^{c}}{\max}|W_{\Phi}(t,s)|]$ in (22), we need to extend the classical Gaussian concentration inequality for Lipschitz functions of real-valued Gaussian elements [23], Dudley’s theorem, and the Borell-TIS inequality [45, 46], to account for complex-valued Gaussian elements arising from the AWT of $\Phi$. To keep the focus on the discussion of ridge deviation, these extensions are provided in Section 5.4.
@@
The proof of Theorem 4 is presented in Section 7.8.
Remark 3 (Interpretation of Theorem 4).
The bound in Theorem 4 is expressed in terms of the threshold $\varepsilon$ and the following ratios:
$\frac{\mu_{1,m}}{L_{m}},\frac{\mu_{2,m}}{L_{m}},\frac{\mu_{3,m}^{2}}{L_{m}},\frac{\mu_{4,m}^{2}}{L_{m}},\frac{\sigma_{m,1}^{2}}{L^{2}_{m}},\frac{\sigma_{m,2}^{2}}{L^{2}_{m}},\frac{\sigma_{m,3}^{2}}{L_{m}},\frac{\sigma_{m,4}^{2}}{L_{m}}.$
To highlight that these ratios become small when the noise strength is low relative to the clean signal, that is, when the signal-to-noise ratio (SNR) is high, we consider the single-component case $M=1$ for brevity. For notational simplicity, below we write $A$, $L$, $\mu_{1},\ldots,\mu_{4}$, and $\sigma_{1},\ldots,\sigma_{4}$ in place of $A_{1}$, $L_{1}$, $\mu_{1,1},\ldots,\mu_{1,4}$, and $\sigma_{1,1},\ldots,\sigma_{1,4}$.
  • 1.
    $L$: A larger value of the second derivative of $|W_{f}(t,\cdot)|^{2}$ with respect to the scale variable at $s_{f}(t)$ indicates that $|W_{f}(t,\cdot)|$ exhibits a sharper peak at $s_{f}(t)$. This can be associated with a rapid change in the spectral content of the signal $f$. Hence, when $L$ is larger, the ridge point of $f$ is less likely to be disturbed. Due to the slowly varying of the magnitude function $A$ and hence the approximation
    $S_{f}(t,s)\approx A^{2}(t)S_{f/A}(t,s),$ (30)
    $L$ is approximately proportional to $A^{2}(t)$. In this case, a larger $A$ reduces the possibility of ridge point disturbance.
  • 2.
    $\mu_{1}$, $\mu_{2}$, $\mu_{3}$, $\mu_{4}$: Under the approximation (30),
    $\frac{6}{L}\max\left\{\mu_{1},\mu_{2},\mu_{3}^{2},\mu_{4}^{2}\right\}\approx\max\left\{\mathcal{O}\left(\frac{\sqrt{\textup{Var}(\Phi)}}{A}\right),\mathcal{O}\left(\frac{\textup{Var}(\Phi)}{A^{2}}\right)\right\}.$ (31)
  • 3.
    For any fixed $\varepsilon$ satisfying (28), (31) shows that the exponents
  • 4.
    and
    This also implies that, for any $\varepsilon$ satisfying (28), the probability that $s_{Y}(t)$ deviates from $s_{f}(t)$ by more than $\varepsilon$ decreases exponentially fast when the ratio $A/\sqrt{\textup{Var}(\Phi)}$ increases.
Remark 4.
The exponential bounds in Theorems 3 and 4 are derived using a generalized Borell-TIS inequality (Lemma 4) for the modulus of Gaussian processes. Since this generalized Borell-TIS inequality neither requires stationarity nor imposes specific structural assumptions on the underlying random process, it is applicable to the nonstationary random processes arising from the modulus of the AWT of an adaptive harmonic signal corrupted by Gaussian noise. On the other hand, the Borell-TIS inequality does not account for the dependence structure induced by the AWT. As a consequence of this generality, the constants in Theorems 3 and 4 are conservative.
@@ 5.3 Numerical simulations
The result is shown in Figure 6. The left panel displays a scatter plot of all $10^{4}$ trials, showing the relationship between the SNR (in dB) and $\triangle$ or $\widetilde{\triangle}$. The right panel presents a box plot summarizing the ridge deviations within different SNR intervals. The result shows that when the SNR is relatively high (resp. low), the ridge point $s_{Y,m}(t)$ of the noisy signal $Y$ is unlikely (resp. likely) to deviate too far from the ridge point $s_{f,m}(t)$ of the clean signal $f$. This numerical result echoes results shown in Theorem 4, providing a bound on the probability that the deviation $|s_{Y,m}(t)-s_{f,m}(t)|$ exceeds a given threshold.
@@ 5.4 Dudley’s theorem and the Borell-TIS inequality for AWT
To derive Theorems 3 and 4, a key challenge is estimating the tail probabilities of the following random variables arising from the AWT of the Gaussian process $\Phi$:
@@
Our estimation method is inspired by the Borell-TIS inequality [45, 46], which provides a tail probability estimate for the supremum of a real-valued non-stationary Gaussian process over an index set. In the proof of the Borell-TIS inequality, the Gaussian concentration inequality for Lipschitz functions of real-valued Gaussian random vectors [23] plays an important role. Lemma 1 below generalizes the concentration inequality in [23] to accommodate complex-valued and circularly symmetric Gaussian random vectors, whose proof is provided in Section 7.9.
@@
Note that in cases (33) and (34), the supremum is taken over the continuous index sets $I$ and $I^{c}$, respectively. In order to apply the dominated convergence theorem to extend the result in Lemma 1 to the continuous index set case, we need to ensure the convergence of the expected value of the supremum. Details of this extension are provided in the proof of Theorem 3. In [14, 45], Dudley’s theorem provides an upper bound for the expected maximum of a centered real-valued Gaussian field over a compact set. Because the AWT of the Gaussian process $\Phi$ is a complex-valued random process and the index set $I^{c}$ for case (33) is noncompact, we present a complex-valued version of Dudley’s theorem for the expectation of the supremum of the process $|W_{\Phi}(t,\cdot)|$ over the noncompact set $0,\infty)$ in Lemma 3. The statement and proof of Lemma 3 rely on estimates of $\mathbb{E}[|W_{\Phi}(0,s)|^{2}]$ for large scale $s$ and the distance function $d_{W_{\Phi}}$ as follows.
@@
In [14, 45], the expectation of the maximum of a centered real-valued Gaussian field over a compact set is bounded above by the Dudley entropy integral. Lemma 3 extends this bound, providing an exact upper bound for the expectation of the maximum of the modulus of a circularly symmetric and nonstationary complex Gaussian process over $(0,\infty)$. The upper bound given in (39) increases as the long-memory parameter $H$ and the decay rate parameter $\gamma$ of the spectral density function of $\Phi$ decrease to zero. This relationship echoes Slepian’s inequality and the Sudakov-Fernique inequality [46, Chapter 7], which state that, for two real-valued Gaussian processes with the same mean, the process with higher correlation exhibits a greater expected maximum. The proof of Lemma 3 is provided in Section 7.11. Finally, to estimate the tail probabilities of the random variables in (35), we introduce the complex-valued version of the Borell-TIS inequality as follows.
@@
The proof of Lemma 4 is provided in Section 7.12. Last but not least, our work bridges and extends classical maximum theorems and maximum inequalities to address a key question in signal processing: how noise perturbs the ridge curves of adaptive harmonic signals that encode the IF information.